Evaluate the ∫e^t dt/cuberoot (1-e^t)


1
Expert's answer
2021-02-24T12:27:11-0500

Solution

Substitution x = et :

"\int \frac{e^t dt}{\sqrt[3]{1-e^t}}=\int\frac{dx}{\sqrt[3]{1-x}}=\int(1-x)^{-1/3}dx=
-\frac{3}{2}(1-x)^{2/3}+C=-\frac{3}{2}(1-e^t)^{2/3}+C"

Answer

"\int \frac{e^t dt}{\sqrt[3]{1-e^t}}=-\frac{3}{2}(1-e^t)^{2/3}+C"